$SPAD/src/input magma.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool magma.output
)set message test on
)set message auto off
)clear all
--S 1 of 22
x:Symbol :='x
--R
--R
--R   (1)  x
--R                                                                 Type: Symbol
--E 1
--S 2 of 22
y:Symbol :='y
--R
--R
--R   (2)  y
--R                                                                 Type: Symbol
--E 2
--S 3 of 22
z:Symbol :='z
--R
--R
--R   (3)  z
--R                                                                 Type: Symbol
--E 3
--S 4 of 22
word := OrderedFreeMonoid(Symbol)
--R
--R
--R   (4)  OrderedFreeMonoid(Symbol)
--R                                                                 Type: Domain
--E 4
--S 5 of 22
tree := Magma(Symbol)
--R

--R
--R   (5)  Magma(Symbol)
--R                                                                 Type: Domain
--E 5
--S 6 of 22
a:tree := x*x
--R
--R
--R   (6)  [x,x]
--R                                                          Type: Magma(Symbol)
--E 6
--S 7 of 22
b:tree := y*y
--R
--R
--R   (7)  [y,y]
--R                                                          Type: Magma(Symbol)
--E 7
--S 8 of 22
c:tree := a*b
--R
--R
--R   (8)  [[x,x],[y,y]]
--R                                                          Type: Magma(Symbol)
--E 8
--S 9 of 22
left c
--R
--R
--R   (9)  [x,x]
--R                                                          Type: Magma(Symbol)
--E 9
--S 10 of 22
right c
--R
--R
--R   (10)  [y,y]
--R                                                          Type: Magma(Symbol)
--E 10
--S 11 of 22
length c
--R
--R
--R   (11)  4

--R                                                        Type: PositiveInteger
--E 11
--S 12 of 22
c::word
--R
--R
--R          2 2
--R   (12)  x y
--R                                              Type: OrderedFreeMonoid(Symbol)
--E 12
--S 13 of 22
a < b
--R
--R
--R   (13)  true
--R                                                                Type: Boolean
--E 13
--S 14 of 22
a < c
--R
--R
--R   (14)  true
--R                                                                Type: Boolean
--E 14
--S 15 of 22
b < c
--R
--R
--R   (15)  true
--R                                                                Type: Boolean
--E 15
--S 16 of 22
first c
--R
--R
--R   (16)  x
--R                                                                 Type: Symbol
--E 16
--S 17 of 22
rest c
--R
--R
--R   (17)  [x,[y,y]]
--R                                                          Type: Magma(Symbol)

--E 17
--S 18 of 22
rest rest c
--R
--R
--R   (18)  [y,y]
--R                                                          Type: Magma(Symbol)
--E 18
--S 19 of 22
ax:tree := a*x
--R
--R
--R   (19)  [[x,x],x]
--R                                                          Type: Magma(Symbol)
--E 19
--S 20 of 22
xa:tree := x*a
--R
--R
--R   (20)  [x,[x,x]]
--R                                                          Type: Magma(Symbol)
--E 20
--S 21 of 22
xa < ax
--R
--R
--R   (21)  true
--R                                                                Type: Boolean
--E 21
--S 22 of 22
lexico(xa,ax)
--R
--R
--R   (22)  false
--R                                                                Type: Boolean
--E 22
)spool
)lisp (bye)

———-

References^{[1] nothing}

7